OpenStudy (anonymous):

f(x)=((1)/(5))(x+9)^2+8 find the vertex

OpenStudy (anonymous):

Equation \[f(x)=1/5(x+9)^{2}+8\]=\[1/5x^2+18/5x+121/5\] = \[ax^2+bx+c\] Therefore, a=1/5 b=18/5 c=121/5 And you can find the X-Axis of the Vertex x=-b/2a=-(18/5)/(2/5)=-9 VERTEX=(-9,8)

OpenStudy (anonymous):

Thanks for explaining :)

OpenStudy (anonymous):